# NEDS-SLAM: A Novel Neural Explicit Dense Semantic SLAM Framework using 3D Gaussian Splatting

### v1

@@ Abstract
1 Introduction
@@
2 Related Work
Gaussian Splatting based SLAM Semantic information is paramount for SLAM systems, as it enables scene understanding beyond geometric reconstruction alone. The ability to perceive and model semantics represents a crucial requirement for enabling applications of SLAM in robotics, virtual or augmented reality, and other domains that require understanding of the environment [17],[5]. Real-time dense semantic SLAM systems face the challenge of effectively fusing semantic information into underlying 3D geometric representations of the environment. Traditional approaches using voxels, point clouds, and signed distance fields to encode object labels [6],[13]. However, voxel- and point cloud-based approaches struggle with reconstruction speed and high-fidelity model acquisition. Meanwhile, signed distance field representations incur high memory usage that does not scale well to large-scale environments. There remains a need for more efficient and expressive 3D semantic modeling techniques suitable for real-time dense SLAM.
Gaussian Splatting based SLAM 3D Gaussian representations have emerged as a promising approach for 3D scene modelling using a set of 3D Gaussians, each characterized by parameters such as position, anisotropic covariance, opacity, and color [9]. While existing GS-based SLAM methods have primarily focused on RGB reconstruction, exploring end-to-end system architectures, optimization of GS parameters, and accurate camera pose tracking through differentiable rendering, less attention has been paid to semantic reconstruction [8],[21],[12],[24] The few semantic GS SLAM approaches proposed to date have simply encoded ground truth semantic color labels directly as a second color channel of the GS parameters [11], without explicit modeling of semantic information or inference. There is clear potential for more sophisticated integration of semantics within the GS framework. The present work conducts a more in-depth exploration of semantic GS SLAM, aiming to simultaneously improve the robustness and reconstruction fidelity of GS-based SLAM systems through more sophisticated modeling and inference of semantic information within the GS representation.
@@ 2 Related Work
3 Methodology
3.1 Scene Representation and Semantic embedding
@@
3.2 Adaptive 3D Gaussian Expansion Mapping
3.2.1 Spatially Consistent Semantic Estimation
@@ 3.2.2 Updating 3D Gaussians
After the modeling process discussed in Section 3.1, the modeled scene contains three feature channels: spatial position, surface color, and potential semantics. The spatial position and surface color are directly obtained from the RGBD data stream. Meanwhile, the fusion of semantic encoding is supervised by the mask output from a pretrained segmentation model.
@@
3.2.3 Pruning 3D Gaussians
@@
The virtual multi-view consistency check method takes advantage of the fast rendering capabilities of the Gaussian Splatting model, enabling the marking of GS points that significantly deviate from the object surface. In subsequent optimization processes, the involvement of outlier GS in the scene is diminished by degrading their opacity. Consistent with [9], GS with near-zero opacity or excessive radius are removed in the mapping process.
3.2.4 Camera tracking
@@
Table 1: Quantitative comparison of map reconstruction and localization accuracy between our proposed NEDS-SLAM and other NeRF-based dense SLAM methods, averaged over 8 scenes from the Replica dataset.
@@
Table 2: We quantitatively compare our proposed NEDS-SLAM method against other radiance field-based SLAM on the ScanNet dataset, reporting the tracking metric of camera pose RMSE measured in centimeters.
4 Experiment
4.1 Experimental Setup
@@
Table 3: Detailed test results of NEDS-SLAM across 8 scenes from the Replica dataset.
@@
Table 4: An ablation study of the NEDS-SLAM, where ’VV’ represents the virtual view pruning method, ’FF’ represents the spatial-consistent semantic fusion method, and ’SL’ represents the utilizing of semantic loss during training.
4.2 Mapping and Tracking result
Quantitative measures of reconstruction quality using the Replica dataset are presented in Table 1. Our method demonstrates competitive performance when compared to other approaches.
4.3 Semantic Mapping result
The NEDS-SLAM, built upon the foundation of 3DGS, achieves accurate camera localization and semantic reconstruction simultaneously. Table. 4 provides a comparison between our method and other neural Implicit approaches in terms of semantic reconstruction performance. Due to the precise representation of object edges offered by the Gaussian radiance field, methods based on neural explicit approaches bring about significant improvements in semantic reconstruction.
@@ 4.3 Semantic Mapping result
Methods mIoU[%] $\uparrow$ Room0 Room1 Office0
@@
Table 5: A quantitative comparison of our method with existing semantic NeRF-based SLAM approaches on the Replica dataset,,For consistency and fair comparison with other methods, we utilize the ground-truth semantic labels from the replica dataset instead of the predicted semantic labels from pre-trained models
@@
4.4 Ablation Study
Ablation experiments were designed for the Feature Fusion module and Pruning GS module to test the effectiveness of NEDS-SLAM for semantic reconstruction. As shown in Table. 4, following SGS-SLAM’s approach, we directly incorporated semantic parameters into the GS by calling a pre-trained M2F segmentation model on each RGB frame. The segmented pixel class labels were color-coded and embedded into the GS parameters, resulting in the reconstruction effect displayed in the fourth column of Figure. 4, corresponding to the first row of Table. 4. For the Replica Room0 dataset, the M2F model achieved a semantic segmentation mIoU of 52.4. Employing this method for semantic SLAM resulted in an average mIoU of 26.52, which was used as the baseline approach. After incorporating the virtual view pruning method, outlier GS points that affected reconstruction quality were effectively removed. This led to a 5% increase in PSNR, slightly improved camera pose tracking accuracy, while maintaining the quality of semantic reconstruction. Building upon the baseline, the incorporation of a semantic loss along with depth and RGB losses, weighted at 0.5, 1.0, and 1.0 respectively, resulted in a 14% improvement in semantic reconstruction performance. The addition of the spatial-consistent semantic fusion method, which combined the semantic features from the DinoV2 ViT-14 model with the appearance features from the DepthAnything model, led to nearly a 50% improvement in semantic reconstruction performance. A lightweight encoder-decoder compressed the high-dimensional fused features into a low-dimensional representation, which was then embedded into the GS parameters.

### v2

@@ Abstract
Index Terms:
3D Gaussian Splatting; Dense Semantic Mapping; Neural SLAM; 3D Reconstruction.
I Introduction
@@
II Related Work
II-A Traditional approaches to dense semantic SLAM
Semantic information is paramount for SLAM systems, as it enables scene understanding beyond geometric reconstruction alone. The ability to perceive and model semantics represents a crucial requirement for enabling applications of SLAM in robotics, virtual or augmented reality, and other domains that require understanding of the environment [6],[7]. Real-time dense semantic SLAM systems face the challenge of effectively fusing semantic information into underlying 3D geometric representations of the environment. Traditional approaches using voxels, point clouds, and signed distance fields to encode object labels [8],[9]. However, voxel- and point cloud-based approaches struggle with reconstruction speed and high-fidelity model acquisition. Meanwhile, signed distance field representations incur high memory usage that does not scale well to large-scale environments. There remains a need for more efficient and expressive 3D semantic modeling techniques suitable for real-time dense SLAM.
II-B Gaussian Splatting based SLAM
3D Gaussian representations have emerged as a promising approach for 3D scene modelling using a set of 3D Gaussians, each characterized by parameters such as position, anisotropic covariance, opacity, and color [10]. While existing GS-based SLAM methods have primarily focused on RGB reconstruction, exploring end-to-end system architectures, optimization of GS parameters, and accurate camera pose tracking through differentiable rendering, less attention has been paid to semantic reconstruction [11],[12],[13],[14] The few semantic GS SLAM approaches proposed to date have simply encoded ground truth semantic color labels directly as a second color channel of the GS parameters [4], without explicit modeling of semantic information or inference. There is clear potential for more sophisticated integration of semantics within the GS framework. The present work conducts a more in-depth exploration of semantic GS SLAM, aiming to simultaneously improve the robustness and reconstruction fidelity of GS-based SLAM systems through more sophisticated modeling and inference of semantic information within the GS representation.
@@ II-B Gaussian Splatting based SLAM
III Methodology
III-A Scene Representation and Semantic embedding
@@
III-B Adaptive 3D Gaussian Expansion Mapping
III-B1 Spatially Consistent Feature Fusion
@@ III-B2 Updating 3D Gaussians
After the modeling process discussed in Section III-A, the modeled scene contains three feature channels: spatial position, surface color, and potential semantics. The spatial position and surface color are directly obtained from the RGBD data stream. Meanwhile, the fusion of semantic encoding is supervised by the mask output from a pretrained segmentation model.
@@
III-B3 Vitrual Camera Pruning 3D Gaussians
@@
Figure 4: Rendered virtual camera views on the ScanNet dataset. he middle images provide a zoomed-in illustration of the effectiveness of Virtual Camera Pruning, where ’vv’ denotes virtual camera view. Eliminating outlier Gaussians not only improves rendering quality but also reduces the storage footprint of the map representation.
@@
The virtual multi-view consistency check method takes advantage of the fast rendering capabilities of the Gaussian Splatting model, enabling the marking of GS points that significantly deviate from the object surface. In subsequent optimization processes, the involvement of outlier GS in the scene is diminished by degrading their opacity. Consistent with [10], GS with near-zero opacity or excessive radius are removed in the mapping process. FSGS [22] used a similar method to improve the quality of images synthesized from novel views. Specifically, FSGS uses the new viewpoint to guide the optimization of grown Gaussians towards a reasonable geometry. Meanwhile, the proposed method in this paper uses multiple synthesized images to eliminate inaccurate Gaussians. As illustrated in Figure. 4, we render virtual views and further optimize the GS distribution only for keyframes. The specific approach for generating virtual views is not fixed. Although Gaussian splatting enables extremely fast virtual view synthesis (nearly 300 FPS), introducing too many viewpoints can compromise the system’s real-time performance. We choose four virtual views along the up, down, left, and right directions, which achieves a desirable balance between effectiveness and efficiency.
III-B4 Camera tracking
@@
Table I: Quantitative comparison of map reconstruction and localization accuracy between our proposed NEDS-SLAM and other NeRF-based dense SLAM methods, averaged over 8 scenes from the Replica dataset.
@@
Table II: We quantitatively compare our proposed NEDS-SLAM method against other radiance field-based SLAM on the ScanNet dataset, reporting the tracking metric of camera pose RMSE measured in centimeters.
IV Experiment
IV-A Experimental Setup
@@
Implementation Details. We conducted experiments using a single NVIDIA 4090 GPU, validating on the REPLICA dataset with the mapping iteration set to 40, tracking iteration set to 60, and SCFF iteration set to 50. After obtaining 384 feature channels through the DINO model, we derived 64-dimensional fused features by applying 2D convolutions separately to the Spatial Features. Finally, we obtained three-dimensional features by passing them through an encoder and embedding them into the GS parameters. We use a learning rate of 0.005 and 0.001 respectively for all learnable parameters on Replica and ScanNet datasets. For camera poses, we only employ a learning rate of 0.0005 in tracking.
@@
Table III: Detailed test results of NEDS-SLAM across 8 scenes from the Replica dataset.
@@
Table IV: An ablation study of the NEDS-SLAM, where ’VV’ represents the virtual view pruning method, ’FF’ represents the spatial-consistent semantic fusion method, and ’SL’ represents the utilizing of semantic loss during training.
IV-B Experiment result
Quantitative measures of reconstruction quality using the Replica dataset are presented in Table I. Our method demonstrates competitive performance when compared to other approaches.
The NEDS-SLAM, built upon the foundation of 3DGS, achieves accurate camera localization and semantic reconstruction simultaneously. Table. IV provides a comparison between our method and other neural Implicit approaches in terms of semantic reconstruction performance. Due to the precise representation of object edges offered by the Gaussian radiance field, methods based on neural explicit approaches bring about significant improvements in semantic reconstruction.
@@ IV-B Experiment result
Methods AVG.mIoU[%] $\uparrow$ Room0 Room1 Office0
@@
Table V: A quantitative comparison of our method with existing semantic NeRF-based SLAM approaches on the Replica dataset,For consistency and fair comparison with other methods, we utilize the ground-truth semantic labels from the replica dataset instead of the predicted semantic labels from pre-trained models
@@
Figure 6: The validation results on the Scannet scene0000_00 dataset. The first row indicates the RGB reconstruction results of NEDS-SLAM, the second row indicates the semantic features predicted by M2F, the third row is the semantic reconstruction results without the Spatially Consistent Feature Fusion (SCFF) module, and the fourth row is the results with the SCFF module.
IV-C Ablation Study
Ablation experiments were designed for the Feature Fusion module and Pruning GS module to test the effectiveness of NEDS-SLAM for semantic reconstruction. As shown in Table. IV, following SGS-SLAM’s approach, we directly incorporated semantic parameters into the GS by calling a pre-trained M2F segmentation model on each RGB frame. The segmented pixel class labels were color-coded and embedded into the GS parameters, resulting in the reconstruction effect displayed in the fourth column of Figure. 5, corresponding to the first row of Table. IV. For the Replica Room0 dataset, the M2F model achieved a semantic segmentation mIoU of 52.4. Employing this method for semantic SLAM resulted in an average mIoU of 26.52, which was used as the baseline approach. After incorporating the virtual view pruning method, outlier GS points that affected reconstruction quality were effectively removed. This led to a 5% increase in PSNR, slightly improved camera pose tracking accuracy, while maintaining the quality of semantic reconstruction. Building upon the baseline, the incorporation of a semantic loss along with depth and RGB losses, weighted at 0.5, 1.0, and 1.0 respectively, resulted in a 14% improvement in semantic reconstruction performance. The addition of the spatial-consistent semantic fusion method, which combined the semantic features from the DinoV2 ViT-14 model with the appearance features from the DepthAnything model, led to nearly a 50% improvement in semantic reconstruction performance. A lightweight encoder-decoder compressed the high-dimensional fused features into a low-dimensional representation, which was then embedded into the GS parameters.
As can be seen in Figure. 6, the semantic features calculated by the M2F model were inconsistent (such as the partitions and books on the table). After processing with the SCFF module, the inconsistencies were resolved and NEDS-SLAM output a more complete semantic reconstruction.